# Center of distances of ultrametric spaces generated by labeled trees

### v1

@@
Abstract. The center of distances of a metric space $(X,d)$ is the set $C(X)$ of all $t\in\mathbb{R}^{+}$ for which the equation $d(x,p)=t$ has a solution for each $p\in X$. We prove that the equalities $C(X)=\{0\}$ or $C(X)=\{\operatorname{diam}X,0\}$ hold if $(X,d)$ is an ultrametric space generated by labeled trees. The necessary and sufficient conditions under which $\operatorname{diam}X\in C(X)$ are found.
@@ 1 Introduction
It was shown in [10] that the compact ultrametric spaces generated by labeled star graphs are the completions of totally bounded ultrametric space generated by decreasingly labeled rays. The labeled trees generating separable ultrametric spaces are characterized, up to isomorphism, in [15].
In the present paper we investigate the class $\bf UT$} of all ultrametric spaces which are isometric to spaces generated by labeled trees. The main goal of the paper is the characterization of the center of distances $C(X)$ of $\bf UT$}-spaces $(X,d)$.
@@
Theorem 3 of Section 3 shows that the center of distances $C(X)$ of arbitrary $\bf UT$}-space $(X,d)$ satisfies either the equality
@@
Theorem 7 of Section 5 shows that equality (1) holds for $\bf UT$}-space $(X,d)$ if and only if the diametrical graph of $(X,d)$ has a spanning star subgraph. Corollary 4 of Theorem 8 of Section 5 shows that equality (2) holds for $\bf UT$}-space $(X,d)$ with $|X|\geq 2$ if and only if $(X,d)$ is weakly similar to an unbounded ultrametric space.
@@ 2 Preliminaries
Following [7, p. 46] we say that a metric $d:X\times X\to\mathbb{R}^{+}$ is equidistant if there is a constant $k>0$ such that
@@
Remark 1.
Some properties of the center of distances are described in [2, 3, 24, 25].
@@
The following concept is well-known for finite graphs (see, for example, [8, p. 17]).
@@
Our next definition is a modification of Definition 2.1 from [26].
@@
Corollary 1.
Let $(X,d)$ be an ultrametric space with $|X|\geq 2$, let $D(X)$ be the distance set of $(X,d)$, and let $C(X)$ be the center of distances of $(X,d)$. If $\operatorname{diam}X$ belongs to $D(X)$,
$\operatorname{diam}X\in D(X),$ (11)
then $\operatorname{diam}X$ also belongs to $C(X)$,
$\operatorname{diam}X\in C(X).$ (12)
Proof.
If (11) holds, then the diametrical graph $G_{X}$ is non-empty by Remark 2. Hence $G_{X}$ is complete multipartite by Theorem 1. Now Definitions 3 and 5 imply (12). ∎
@@
with $k\geqslant 2$. We will write $P=(x_{1},\ldots,x_{k})$ or $P=P_{x_{1},x_{k}}$ if $P$ is a path satisfying (13) and said that $P$ is a path joining $x_{1}$ and $x_{k}$. A graph $G$ is connected if for every two distinct vertices of $G$ there is a path $P\subseteq G$ joining these vertices.
@@
The last equality and formula (18) show that $l^{*}\colon V(T)\to\mathbb{R}^{+}$ is a non-degenerate labeling of $V(T)$ because $l\colon V(T)\to\mathbb{R}^{+}$ is non-degenarate. Thus $(V(T),d_{l^{*}})$ is an ultrametric space by Theorem 2.
@@ 3 Center of distances of UT-spaces
The ultrametric $d^{+}$ on $\mathbb{R}^{+}$ was introduced by Delhommé, Laflamme, Pouzet, and Sauer in [6]. Some results related to the ultrametric space $(\mathbb{R}^{+},d^{+})$ can be found in [20, 12, 22, 21, 23].
@@ 4 Center of distances and centered spheres
The next our goal is to describe some conditions under which every open ball of a space $(X,d)\in{\bf UT}$ is a centered sphere in this space.
@@
Similarly (44) we have the inclusion
@@
Let us consider the ultrametric space $(X,d)$ depicted by Figure 1. Then $(X,d)\in{\bf UT}$ and the set $\{x_{1},x_{2}\}$ is a centered sphere in itself but any centered sphere in $(X,d)$ is either a singleton or coincides with one of sets $\{x_{3},x_{4}\}$, $\{x_{1},x_{2},x_{3}\}$, $\{x_{1},x_{2},x_{4}\}$, $\{x_{1},x_{2},x_{3},x_{4}\}$.
@@
for each $p\in X$. Then the following statements are equivalent:
  1. (i)
    The inclusion
@@
    holds.
  2. (ii)
    The set $D_{p}(X)\cap[0,r)$ has the greatest element for all $p\in X$ and $r\in(0,\infty)$.
@@
$(i)\Rightarrow(ii)$. Let inclusion (50) hold, let $p$ be an arbitrary point of $X$, and let $r\in(0,\infty)$. We must show that the set $D_{p}(X)\cap[0,r)$ has the greatest element. This element obviously exists if $|D_{p}(X)\cap[0,r)|=1$.
Let us now assume that $|D_{p}(X)\cap[0,r)|\geq 2$ hold and consider the open ball
@@
Inclusion (49) implies that $B$ is a centered sphere in $(X,d)$. By Lemma 3 the ultrametric space $(B,d|_{B\times B})$ is an $\bf UT$}-space because $(X,d)\in{\bf UT}$ by condition of the theorem. Now, using Theorem 4 for the space $(B,d|_{B\times B})$, we obtain
$\operatorname{diam}B\in C(B)\subseteq D(B),$
where $C(B)$ and $D(B)$ are the center of distances and, respectively, the distance set of $(B,d|_{B\times B})$. Consequently, the diametrical graph $G_{B}$ of $(B,d|_{B\times B})$ is non-empty. Theorem 1 implies that $G_{B}$ is a complete multipartite graph. Hence, by Definition 5, there is a point $x\in B$ such that
@@
$(ii)\Rightarrow(i)$. Let statement $(ii)$ hold. Let us consider an arbitrary open ball $B^{1}\in{\bf B}_{X}$ with a center $p_{1}\in X$ and a radius $r_{1}>0$,
@@
holds. Let $D_{p_{1}}$ be the set defined by (49) with $p=p_{1}$, where $p_{1}$ is the center of the ball $B^{1}$. By statement $(ii)$ the set $D_{p_{1}}\cap[0,r_{1})$ has the greatest element $R_{1}$, where $r_{1}$ is the radius of the ball $B^{1}$. Thus we have
$R_{1}\in D_{p_{1}}\cap[0,r_{1})\subseteq D(B^{1}).$
The strong triangle inequality and (52) imply $R_{1}=\operatorname{diam}B^{1}$. Thus (54) holds and, consequently, we have
@@
by Theorems 3 and 4. Using Lemma 4 we may rewrite (55) as
@@
Thus (56) holds for each $B^{1}\in{\bf B}_{X}$. Inclusion (50) follows. The proof is completed. ∎
@@
Let $(X,d)$ be isometric to $({\mathbb{Q}},d_{p})$, where $d_{p}$ is the $p$-adic ultrametric defined in Example 1. Then using formula (7), it is easy to see that statement $(ii)$ of Theorem 5 is valid. Moreover, since the space $({\mathbb{Q}},d_{p})$ does not contain isolated points, there are no open balls in $({\mathbb{Q}},d_{p})$ that are centered spheres in this space. Thus the inclusion ${\bf B}_{X}\subseteq{\bf Cs}_{X}$ is false.
The next theorem is a direct consequence of Theorem 4.9 from [11].
@@
Let $(X,d)$ be a equidistant metric space. Then each non-empty subset $A$ of $X$ is a centered sphere in $(A,d|_{A\times A})$ but any centered sphere in $(X,d)$ is either a singleton or coincides with the set $X$.
@@ 5 Center of distances, diametrical graphs, and weak similarities
Definition 11.
$d(x,y)=f\left(\rho\left(\Phi(x),\Phi(y)\right)\right)$
for all $x,y\in X$.
The next theorem is a reformulation of Theorem 3.22 from [5].
Theorem 8.
  1. (i)
    $(X,d)$ is weakly similar to an unbounded ultrametric space.
  2. (ii)
    The distance set $D(X)$ does not contain $\operatorname{diam}X$,
    $\operatorname{diam}X\notin D(X).$
The last theorem and Theorem 3 give us the following corollary.
@@
It follows directly from Definition 10, that every centered sphere in the space $(X,d)$ is a closed ball of this space,
${\bf Cs}_{X}\subseteq\overline{\bf B}_{X}.$
@@
for each $p\in X$. Then the following statements are equivalent:
  1. (i)
    The inclusion
    $\overline{{\bf B}}_{X}\subseteq{\bf Cs}_{X}$
    holds.
  2. (ii)
    The set $D_{p}(X)\cap[0,r]$ has the greatest element for all $p\in X$ and $r\in\mathbb{R}^{+}$.
@@
Then direct calculations show that every non-empty subset of $X_{3}$ is a centered sphere in $(X_{3},d)$.
@@
The following conjecture was motivated by properties of $p$-adic ultrametric with $p=2$ (see Example 1).
@@
$C(Y)=1+\lfloor\log_{2}(n)\rfloor.$
Acknowledgement

### v2

@@
Abstract. The center of distances of a metric space $(X,d)$ is the set $C(X)$ of all $t\in\mathbb{R}^{+}$ for which the equation $d(x,p)=t$ has a solution for each $p\in X$. We prove that the equalities $C(X)=\{0\}$ or $C(X)=\{0,\operatorname{diam}X\}$ hold if $(X,d)$ is an ultrametric space generated by labeled trees. The necessary and sufficient conditions under which $\operatorname{diam}X\in C(X)$ are found.
@@ 1 Introduction
It was shown in [8] that the compact ultrametric spaces generated by labeled star graphs are the completions of totally bounded ultrametric space generated by decreasingly labeled rays.
In the present paper we investigate the class $\bf UT$} of all ultrametric spaces which are isometric to spaces generated by labeled trees. The main goal of the paper is the characterization of the center of distances of $\bf UT$}-spaces $(X,d)$. The notion of center of distances was introduced in paper [2] and used for a generalization of John von Neumann Theorem on permutations of two sequences with the same set of cluster points [22]. At the moment the authors are not aware of any results describing the structure of the center of distances in the case of ultrametric spaces.
@@
Theorem 4 of Section 3 shows that the center of distances $C(X)$ of an arbitrary $\bf UT$}-space $(X,d)$ satisfies either the equality
@@
Theorem 8 of Section 5 shows that equality (1) holds for $\bf UT$}-space $(X,d)$ if and only if the diametrical graph of $(X,d)$ has a spanning star subgraph. Corollary 4 of Theorem 1 shows that equality (2) holds for $\bf UT$}-space $(X,d)$ with $|X|\geq 2$ if and only if $(X,d)$ is weakly similar to an unbounded ultrametric space.
The final Section 6 contains some conjectures related to results of the paper.
@@ 2 Preliminaries
Following [5, p. 46] we say that a metric $d:X\times X\to\mathbb{R}^{+}$ is equidistant if there is a constant $k>0$ such that
@@
Definition 4.
$d(x,y)=f\left(\rho\left(\Phi(x),\Phi(y)\right)\right)$
for all $x,y\in X$.
The next theorem is a reformulation of Theorem 3.22 from [3].
Theorem 1.
  1. (i)
    $(X,d)$ is weakly similar to an unbounded ultrametric space.
  2. (ii)
    The distance set $D(X)$ does not contain $\operatorname{diam}X$,
    $\operatorname{diam}X\notin D(X).$
Lemma 1.
Let $A$ be a non-empty subset of an ultrametric space $(X,d)$. Then the equality
$\operatorname{diam}A=\sup\{d(p,a):a\in A\}$ (9)
holds for each $p\in A$.
Proof.
Let $p$ be an arbitrary point of $A$. Then, using the strong triangle inequality, we obtain
$d(x,y)\leq\max\{d(p,x),d(p,y)\}\leq\sup\{d(p,a)\colon a\in A\}.$
Consequently, the inequality
$\operatorname{diam}A\leq\sup\{d(p,a)\colon a\in A\}$
holds. The converse inequality follows directly from (8). Thus equality (9) holds. ∎
Remark 1.
Equality (9) is known, see, for example, equality (1.3) in [15], but the authors cannot give here an exact reference to standard books on ultrametric spaces.
@@
The following concept is well-known for finite graphs (see, for example, [6, p. 17]).
@@
The next concept was introduced in [19] for characterization of finite ultrametric spaces for which the Gomory–Hu inequality [17] turns to equality.
@@
with $k\geqslant 2$. We will write $P=(x_{1},\ldots,x_{k})$ or $P=P_{x_{1},x_{k}}$ if $P$ is a path satisfying (12) and say that $P$ is a path joining $x_{1}$ and $x_{k}$. A graph $G$ is connected if for every two distinct vertices of $G$ there is a path $P\subseteq G$ joining these vertices.
@@
The last equality and formula (17) show that $l^{*}\colon V(T)\to\mathbb{R}^{+}$ is a non-degenerate labeling of $V(T)$ because $l\colon V(T)\to\mathbb{R}^{+}$ is non-degenerate. Thus $(V(T),d_{l^{*}})$ is an ultrametric space by Theorem 3.
@@ 3 Center of distances of UT-spaces
Lemma 1 and Proposition 3 imply the following corollary.
Corollary 1.
Let $(X,d)$ be an ultrametric space, $D(X)$ be the distance set of $(X,d)$ and $C(X)$ be the centers of distances of $(X,d)$. If $\operatorname{diam}X$ belongs to $D(X)$,
$\operatorname{diam}X\in D(X),$
then $\operatorname{diam}X$ also belongs to $C(X)$,
$\operatorname{diam}X\in C(X).$
@@
The ultrametric $d^{+}$ on $\mathbb{R}^{+}$ was introduced by Delhommé, Laflamme, Pouzet, and Sauer in [4].
@@ 4 Center of distances and centered spheres
Our next goal is to describe some conditions under which every open ball of a space $(X,d)\in{\bf UT}$ is a centered sphere in this space.
@@
Similar to (43), we have the inclusion
@@
Let us consider the ultrametric space $(X,d)$ depicted by Figure 1. Then $(X,d)\in{\bf UT}$ and the set $\{x_{1},x_{2}\}$ is a centered sphere in itself but this set is not a centered sphere in $(X,d)$. Indeed, the equality $d(x_{1},x_{2})=2$ and the definition of $(X,d)$ give us the equalities
$\{x\in X\colon d(x,x_{1})=2\}\cup\{x_{1}\}=\{x_{1},x_{2},x_{3},x_{4}\}=\{x\in X\colon d(x,x_{2})=2\}\cup\{x_{2}\}.$
@@
for each $p\in X$. If the inclusion
@@
holds, then the set $D_{p}(X)\cap[0,r)$ has the greatest element for all $p\in X$ and $r\in(0,\infty)$. Moreover, if $(X,d)\in{\bf UT}$ and the set $D_{p}(X)\cap[0,r)$ has the greatest element for all $p\in X$ and $r\in(0,\infty)$, then inclusion (49) holds.
@@
Let inclusion (49) hold, let $p$ be an arbitrary point of $X$, and let $r\in(0,\infty)$. We must show that the set $D_{p}(X)\cap[0,r)$ has the greatest element.
Let us consider the open ball
@@
By inclusion (49) the ball $B$ is a centered sphere in $(X,d)$, i.e., the equality
$B=\{x\in X\colon\ d(x,p_{1})=r_{1}\}\cup\{p_{1}\}$
holds with some $p_{1}\in B$ and $r_{1}\in\mathbb{R}^{+}$. The last equality and the Lemma 1 with $A=B$ and $p=p_{1}$ imply
$\operatorname{diam}B=r_{1}\quad\text{and}\quad\operatorname{diam}B\in D(B),$
where $D(B)$ is the distance set of $(B,d|_{B\times B})$.
Consequently, the diametrical graph $G_{B}$ of $(B,d|_{B\times B})$ is non-empty. Theorem 2 implies that $G_{B}$ is a complete multipartite graph. Hence, by Definition 6, there is a point $x\in B$ such that
@@
Let $(X,d)\in{\bf UT}$ and let the set $D_{p}(X)\cap[0,r)$ contain the greatest element for all $p\in X$ and $r\in(0,\infty)$. Let us consider an arbitrary open ball $B^{1}\in{\bf B}_{X}$ with a center $p_{1}\in X$ and a radius $r_{1}>0$,
@@
holds. Let $D_{p_{1}}(X)$ be the set defined by (48) with $p=p_{1}$, where $p_{1}$ is the center of the ball $B^{1}$. By supposition the set $D_{p_{1}}(X)\cap[0,r_{1})$ has the greatest element $R_{1}$. Thus we have
$R_{1}\in D_{p_{1}}(X)\cap[0,r_{1})\subseteq D(B^{1}).$
Lemma 1 and (51) imply $R_{1}=\operatorname{diam}B^{1}$. Thus (53) holds and, consequently, we have
@@
by Theorems 4 and 5. Using Lemma 5 we may rewrite (54) as
@@
Thus (55) holds for each $B^{1}\in{\bf B}_{X}$. Inclusion (49) follows.
The proof is completed. ∎
@@
Let $(X,d)$ be isometric to $({\mathbb{Q}},d_{p})$, where $d_{p}$ is the $p$-adic ultrametric defined in Example 1 and let
$D_{q}(X):=\{d(q,x)\colon x\in X\}$
for $q\in X$. Then using formula (7), it is easy to see that the set $D_{q}(X)\cap[0,r)$ contains the greatest element for all $q\in X$ and $r\in(0,\infty)$. Since the space $({\mathbb{Q}},d_{p})$ does not contain isolated points, there are no open balls in $({\mathbb{Q}},d_{p})$ that are centered spheres in this space. Thus the inclusion ${\bf B}_{X}\subseteq{\bf Cs}_{X}$ is false. It should be noted that singletons are centered spheres by Proposition 5 but such sets cannot be open balls if the space does not contain isolated points.
It is interesting to compare Theorem 6 with the following result.
@@
Let $(X,d)$ be an equidistant metric space. Then each non-empty subset $A$ of $X$ is a centered sphere in $(A,d|_{A\times A})$ but any centered sphere in $(X,d)$ is either a singleton or coincides with the set $X$.
@@ 5 Center of distances, diametrical graphs, and weak similarities
Theorems 1 and 4 give us the following.
@@
for each $p\in X$. If the inclusion
$\overline{{\bf B}}_{X}\subseteq{\bf Cs}_{X}$ (62)
holds, then the set $D_{p}(X)\cap[0,r]$ has the greatest element for all $p\in X$ and $r\in\mathbb{R}^{+}$. Moreover, if $(X,d)\in{\bf UT}$ and the set $D_{p}(X)\cap[0,r]$ has the greatest element for all $p\in X$ and $r\in\mathbb{R}^{+}$, then inclusion (62) holds.
@@
Then it is easy to show that every non-empty subset of $X_{3}$ is a centered sphere in $(X_{3},d)$.
@@
The following conjecture was motivated by some calculations with $p$-adic ultrametric with $p=2$ (see Example 1).
@@
$|C(Y)|=1+\lfloor\log_{2}(n)\rfloor.$
To see why a binary logarithm bound is expected in Conjecture 3, we only note that the following simple statement is valid.
Let $\{X_{1},X_{2},\dots,X_{k}\}$ be a partition of a finite set $X$ on disjoint subsets $X_{1},X_{2},\dots,X_{k}$ with $k\geq 2$. Then there is a partition $\{X_{1}^{*},X_{2}^{*}\}$ of $X$ such that
$\min\{|X_{1}^{*}|,|X_{2}^{*}|\}\geq\min\{|X_{1}|,\dots,|X_{k}|\}.$
Funding